November 21CMS Service Interruption 3:40am - 7:00am due to network issues.
November 18CMS Service was under heavy load 8:40am - 12:30pm. All websites were slow to respond during this period of load.
October 23CMS Service was manually restarted. All websites stopped serving for approximately 10-15 minutes. Service resumed normal state at 6:00pm.
October 18CMS Service was manually restarted due to approximately 20 websites repeatedly relaunching. The CMS Service restart took much longer than anticipated. All websites stopped serving for approximately 15 minutes. Service resumed normal state at 12:40pm.
October 9Between the hours of 1am and 5am our data centre provider carried out network maintenance to address issues noted from the July outage. During this time frame, there were several short periods where connectivity was interupted.
September 29/30Our CMS hosting services experienced issues that intermittenly made browser connection attempts fail. The timeframe was approximately 11pm - 1:30am.
September 26Our CMS hosting services experienced issues that intermittenly made browser connection attempts fail. The timeframe was approximately 6am till midday.
August 14Our hosting services experienced issues related to a massive number of concurrent connections. Service has been restored. Timeframe was 2:14am - 7:23am
July 19A network issue impacted some services including email delivery from website forms, and some internal http requests. Data centre engineers resolved the issue at approximately 8:00pm.
July 13Data centre engineers implemented a fix for packet loss and slowness over multiple regions in response to the earlier service interuption.
July 13At 9:40am an outage that affected our data centre suppliers operations across Australia/New Zealand caused all network traffic to and from our servers to cease. This was resolved at approximately 11:20am. Data centre engineers are still working on their systems, and there could be further short periods of network interuptions.
July 6Planned Maintenance: Router reboot scheduled for 11:30pm. Services will be interrupted for approximately 5 minutes.
July 3Our Hosting service experienced an issue at the data connection phase. Relevant systems/services were restarted, and normal service resumed. Time frame for this was 7:36pm till 8:10pm. This event is still under investigation.
May 11Our CMS Hosting service experienced an issue at the data connection phase. Relevant systems/services were restarted, and normal service resumed. Time frame for this was 10:50am till 10:57am.
May 4Our hosting services experienced issues related to a massive number of concurrent connections. Service has been restored. Time frame for this was 10:10am till 10:30am.
April 26There was a problem with our data centre's upstream connection provider on the 26th of April 2018, which allowed a DDOS attack to penetrate through and affect our services. Our services were affected for approximately 15 minutes (late afternoon).
April 4PHP Service interuption: A set of very large backup files exposed an mismatch between disk space provisioning and disk space allocation, and as a result the service effectively had zero free space. This was resolved at approximately 10:00am.
October 28Some HTTPS website experienced issues, and would have reported as insecure to website viewers. This was resolved at approximately 10:30am.
October 1Webmail was temporarily unavailable - Service was restored at 1pm
July 20Email Hosting experienced a sustained DDOS attack for approximately 2.5 hours (9:30am - 12:00pm). This resulted in some users being unable to send during this time.
July 19Email Hosting experienced a DDOS attack for just over 1 hour (9:30am - 10:45am). This resulted in some users being unable to send during this time.
July 16CMS hosting had a temporary period of slowness and time outs 6:55am - 7:40am.
July 13Email service temporary interruption may result in some delayed email.
July 11CMS Hosting services impacted by issues related to load. Certain websites were targeted during this time. Our data centre has been assisting us with these issue.
Services were restored at 10:00pm